OpenStudy (anonymous):

Help :/ find all the roots by factoring or using the quadratic formula.. f(x)=x^4+5x^2+4

OpenStudy (accessdenied):

To find the roots, we set f(x) = 0. \(x^4 + 5x^2 4 = 0\) Now, at first glance, this does not appear to be a quadratic equation. However, what happens if we made a substitution like u = x^2 ?

OpenStudy (accessdenied):

oops, that should say \(x^4 + 5x^2 + 4 = 0\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm stuck here...l

OpenStudy (anonymous):

(4x^2) (4x^2-1)

OpenStudy (phi):

with u= x^2 u^2 +5u +4 =0

OpenStudy (accessdenied):

^ Yep. In this form, it becomes a quadratic equation! So, now we are able to either factor or use quadratic formula with this equation.